OpenStudy (anonymous):

V = π*r 2 *h V=π*(16) 2 *24=19301.9inches^ 3   S.A.=2*π*r 2 +2*π*r*h = (2)(pi)(16)(16) + (2)(pi)(16)(24) = 1280πinches^ 2 = 4021.2inches^ 2

OpenStudy (anonymous):

someone help me do this but it not good can someone look it over for me this is the comment i got for it from my Instructor In general, we can use the superscript which is in the editor by the Greek character (such as x²) or you can use ^ such as x^2 to indicate raising to a power. For the area you are correct in that we will use (PI)(r^2)(h) to compute the volume. The height is 24 inches as you noted. But the radius is not 16. For the surface area, it looks like you are using the correct formula, but again, the radius will not be 16.

OpenStudy (anonymous):

\(\text{Volume}_{\text{Cylinder}} = 2 \pi r^2 h\) It's r^2 not r*2 But otherwise yes you're good :-) And ty for remembering the units!

OpenStudy (anonymous):

good job! ^_^

OpenStudy (anonymous):

thank you so that all i need to change in it after that everything is good :)

OpenStudy (anonymous):

or u can just remember that a cylinder is a circle with h

OpenStudy (anonymous):

so formula would be V=Ah

OpenStudy (anonymous):

and area is area of circle

OpenStudy (anonymous):

but she said it not 16

OpenStudy (anonymous):

I'm still lost with this

OpenStudy (anonymous):

where is the original question

OpenStudy (anonymous):

your instructor says everything is fine except the radius. you may be using diameter in place of radius. check once

OpenStudy (anonymous):

This is what I had the frist time

OpenStudy (anonymous):

MODELING - Complementary Angles The difference between the measures of two complementary angles is 16°. Determine the measures of the two angles. x+y=90 x-y= 16 add them: 2x = 106 x = 53 Get y: y = 90-53 = 37 They are 37 degrees and 53 degrees

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is that good

OpenStudy (anonymous):

thats correct. how this is related to volume question ?

OpenStudy (anonymous):

sorry post thr rong one

OpenStudy (anonymous):

V = π*r 2 *h V=π*(16) 2 *24=19301.9inches^ 3   S.A.=2*π*r 2 +2*π*r*h = (2)(pi)(16)(16) + (2)(pi)(16)(24) = 1280πinches^ 2 = 4021.2inches^ 2

OpenStudy (anonymous):

this is the solution you worked out. where is the question ?

OpenStudy (anonymous):

r = 6 inches in the attachment you're using r=16 wrongly

OpenStudy (anonymous):

thats the mistake in radius your instructor is talking about :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

In Exercises 7–16, determine (a) the volume and (b) the surface area of the three-dimensional figure. When appropriate, use the key on your calculator and round your answer to the nearest hundredth

OpenStudy (anonymous):

that is the question that come with it

OpenStudy (anonymous):

it have the pi in it after use the

OpenStudy (anonymous):

replace 16 with 6 in your work. you will be fine

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you have put r=16 in your work replace it to r = 6 :)

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

no problem :D

OpenStudy (anonymous):

so it would look like this V = π*r 2 *h V=π*(6) 2 *24=19301.9inches^ 3   S.A.=2*π*r 2 +2*π*r*h = (2)(pi)(6)(6) + (2)(pi)(6)(24) = 1280πinches^ 2 = 4021.2inches^ 2

OpenStudy (anonymous):

no you need to replace r and do the calculation again

OpenStudy (anonymous):

replace r with 6

OpenStudy (anonymous):

im sorry im just bad at this

OpenStudy (anonymous):

V = π*r^2 *h V=π*(6)^2 *24=2714.3 inches^ 3   S.A.=2*π*r^2 +2*π*r*h = (2)(pi)(6)(6) + (2)(pi)(6)(24) = 1131.0 inches^ 2

OpenStudy (anonymous):

hmm its okay for now. you should learn to use the formula atleast. its not that hard as it seems. pls try once. you can ask ppl here to teach you how to use formula, they will be very glad to teach you the formula rather than do the whole problem for you. hope you understand me :)

OpenStudy (anonymous):

thank you and yes I will do that

OpenStudy (anonymous):

okay :) all the best :DDD